# Activating Distributed Visual Region within LLMs for Efficient and Effective Vision-Language Training and Inference

### v1

@@ 4.2 What is the necessary scale of layers for effective and efficient LVLMs training?
To investigate the necessary scale of this visual region to enable LVLMs to receive visual signals and align with linguistic features, we re-train Bunny-Llama-3-8B-V by updating varying number of layers. We respectively experiment with configurations of 32, 16, 8, 6, 4, 2 and 1 layers, with all selected layers uniformly distributed across all layers  [^2] Specifically, we select all even-numbered layers for the 16-layer configuration; layer 0, 4, 8, 12, 18, 22, 26, 30 for the 8-layer setup; layer 0, 6, 12, 18, 24, 30 for the 6-layer setup; and layer 0, 10, 20, 30 for the 4-layer configuration (Our experiments show that using layer 30 or 31 leads to comparable results). For 2 layers and 1 layers that can not selected uniformly, we choose layers with highest block influence scores, i.e., layer 0 and 31 for 2 layers, and layer 31 for 1 layer.. This selection strategy is based on our finding that sparsely and uniformly distributed layers are the optimal position for effective visual learning.
@@
Model Scale Perception Cognition Avg
OCRVQA DocVQA RefCOCOg TDIUC MMBench GQA ScienceQA TextVQA
@@
32 layers 61.51% 19.46% 49.01% 83.40% 66.67% 62.98% 68.47% 58.19% 58.71%
16 layers 64.01% 20.75% 48.02% 83.47% 64.00% 62.43% 67.53% 58.27% 58.56% (99.74%)
8 layers 62.19% 21.10% 47.71% 83.10% 63.92% 61.60% 68.17% 57.35% 58.14% (99.03%)
6 layers 61.39% 22.84% 46.54% 83.31% 61.77% 61.08% 68.32% 56.19% 57.69% (98.26%)
4 layers 63.28% 21.01% 43.47% 83.14% 60.82% 60.48% 67.97% 54.48% 56.83% (96.80%)
2 layers 54.54% 19.10% 41.90% 81.47% 57.22% 57.38% 65.84% 53.27% 53.84% (91.70%)
1 layer 53.16% 16.96% 33.29% 81.20% 51.89% 55.83% 64.50% 45.51% 50.29% (85.66%)
@@
40 layers 67.60% 25.19% 50.26% 83.61% 68.38% 63.29% 71.64% 60.21% 61.27%
10 layers 65.17% 23.56% 48.27% 83.57% 66.58% 62.01% 70.75% 59.13% 59.88% (97.73%)
9 layers 66.47% 23.65% 49.29% 83.74% 65.61% 62.31% 72.14% 59.71% 60.37% (98.53%)
Table 3: Performance Comparison of LLaVA-1.5 with different model scales tuned with different numbers of layers. Bold numbers represent the best performance in each column. Values in parentheses denotes the percentage relative to the performance achieved by tuning all layers.
@@ 5.1 Generalizability Validation
To validate our findings of visual region beyond Bunny-Llama-3-8B-V, we take LLaVA-1.5-7B and LLaVA-1.5-13B as additional testbeds to assess the generalizability across LVLMs with different LLM backbones and parameter scales. Following the setup in Sec. 4.2, we re-train both models with different number of layers that are sparsely and uniformly distributed within their respective backbones, Vicuna-1.5-7B and Vicuna-1.5-13B. Results presented in Table 3 show that under our visual region positioning strategy, tuning approximately 25% of the layers consistently yield 98% of the full performance. This demonstrates that our approach generalizes effectively across LVLMs.
@@ 5.2 Computational Cost
To demonstrate the efficiency of visual region-based tuning, we report the computational costs associated with tuning different numbers of layers across various models using the LoRA strategy. For fair comparison across setups with different numbers of GPUs (specifically A800 GPUs in this analysis), we compute the product of the number of GPUs and running hours as a measure of computational cost. From Figure 3, Table 2 and Table 3, tuning a visual region comprising up to 25% of layers (8 layers for LLaVA-1.5-7B and Bunny-Llama3-8B-V, 10 layers for LLaVA-1.5-13B) can achieve 98% of full performance while achieving
@@
substantial reductions in computational overhead. Specifically, we reduce training time by 23% for LLaVA models and 13% for Bunny. These results highlight that the effectiveness of visual region-based tuning in training LVLMs efficiently with minimal performance trade-offs.
@@ 5.3 Evaluation of Textual Tasks
As highlighted in Dai et al. (2024); Agrawal et al. (2024) and illustrated in Figure 1, multimodal training risks significant degradation of LLMs’ pre-trained linguistic knowledge and reasoning capabilities. To verify whether training our sparsely and uniformly distributed visual region compromises the model linguistic capacity, we extend our experiments to two text-only question answering datasets, MMLU Hendrycks et al. (2020) and BIG-bench-Hard Suzgun et al. (2022), covering a wide range of topics and fields. We use “Answer with the option’s letter from the given choices directly” and “Please answer this question in a word or phrase” as the prompts for MMLU and BIG-bench-Hard. We calculate the multi-choice accuracy as the evaluation metric, allowing models to provide additional explanations alongside its responses. We adopt a five-shot prompting strategy for MMLU and a zero-shot strategy for BIG-bench-Hard.
Model Version MMLU BIG-bench-hard
@@
Fully-trained (32 layers) 60.27% 30.93%
Partial-trained( 8 layers) 63.36% 31.50%
LLM-Backbone 66.01% 57.93%
@@
Fully-trained (32 layers) 50.52% 26.85%
Partial-trained (8 layers) 50.74% 31.64%
LLM-Backbone 49.78% 29.33%
Table 4: Performance on text-only tasks. The LLm backbones of Bunny-LLaMA3-8B-V and LLaVA-1.5-7B are respectivly LLaMA3-8B and Vicuna-1.5-7B.
As shown in Table 4, fully-trained LVLMs generally exhibit decreased performance on text-only tasks compared to their LLM backbones, particularly on the challenging BIG-bench-Hard dataset. In contrast, our selectively trained LVLMs consistently achieve better performance than fully-trained LVLMs, even sometimes outperform their LLMs backbones. These results support our hypothesis regarding the positioning of visual region, that tuning sparsely and uniformly distributed layers better preserves LLMs’ pre-existing linguistic knowledge and reasoning capabilities, while full training may cause minor disruptions.
@@ 6 Visual Region-Based Layer Pruning
Building on our visual region targeted training, we propose a visual region-based pruning paradigm that selectively prunes less-important layers outside the visual region after training. Specifically, we follow the angular distance based layer importance metric and select 0$\sim$4 layers with the lowest angular distance outside the visual region. We do not evaluate pruning beyond this range as removing additional layers in LVLMs would lead to significant performance collapse. We evaluate this approach on LLaVA-1.5-7B across four datasets: OCRVQA, TextVQA, DocVQA and SciQA. As shown in Figure 4, our paradigm generally maintain higher performance, especially when pruning 3$\sim$4 layers, even though the visual region targeted trained model performs slightly worse than fully trained model without pruning. This result demonstrates that our paradigm effectively minimizes performance degradation compared to pruning in full-layer trained LVLMs.
@@ 7.1 Efficient Training and Inference
Recent research community has witnessed an emergent interest in LLMs Touvron et al. (2023); Chiang et al. (2023) and LVLMs Li et al. (2023); Zhu et al. (2023); Bai et al. (2023); Liu et al. (2024) due to their remarkable ability to interpret and interact with the world via linguistic and visual channels. With the sustainably increased scale of LLMs and LVLMs, training or inference using all model parameters are cost for practical deployment. There are numerous techniques for efficient model training and inference. For instance, quantization reduce the memory footprint of models by decreasing the precision of model weights Dettmers et al. ; Dettmers and Zettlemoyer (2023); Xiao et al. (2023). Low rank adapters enable cost-effective fine-tuning by updating only a small subset of the adapter parameters Hu et al. (2021); Karimi Mahabadi et al. (2021).
@@ 8 Conclusion
Our work primarily focuses on LLaVA-1.5 family and Bunny-LLama3-8B-V to demonstrate the effectiveness and efficiency of our proposed training and inference paradigms for LVLMs. Future work will expand to a broader range of models to further validate the generalizability of our approach. Additionally, we will explore extensions to other modalities such as speech, and investigate the existence of other modality-specific regions to develop more versatile and scalable multimodal models.

### v2

@@ 4.2 What is the necessary scale of layers for effective and efficient LVLMs training?
To investigate the necessary scale of this visual region to enable LVLMs to receive visual signals and align with linguistic features, we re-train Bunny-Llama-3-8B-V by updating varying number of layers. We respectively experiment with configurations of 32, 16, 8, 6, 4, 2 and 1 layers, with all selected layers uniformly distributed across all layers  [^2] Specifically, we select all even-numbered layers for the 16-layer configuration; layer 0, 4, 8, 12, 18, 22, 26, 30 for 8-layer; layer 0, 6, 12, 18, 24, 30 for 6-layer; and layer 0, 10, 20, 30 for 4-layer (Experiments show that layer 30 or 31 yields comparable results, and all odd-numbered selections perform slightly worse). Since 2-layer and 1-layer selection can not be uniform, we have tested various configurations and adopted the best-performing strategy: layer 0 and 31 for 2-layer, and layer 31 for 1-layer based on highest block influence scores.. This selection strategy is based on our finding that sparsely and uniformly distributed layers are the optimal position for effective visual learning.
@@
Model Scale OCRVQA DocVQA RefCOCOg TDIUC MMBench GQA ScienceQA TextVQA MMMU SEED-IMG Avg
@@
32 layers 61.51% 19.46% 49.01% 83.40% 66.67% 62.98% 68.47% 58.19% 35.3% 67.52% 57.25%
16 layers 64.01% 20.75% 48.02% 83.47% 64.00% 62.43% 67.53% 58.27% 35.4% 67.22% 57.11%(99.76%)
8 layers 62.19% 21.10% 47.71% 83.10% 63.92% 61.60% 68.17% 57.35% 34.6% 67.23% 56.70%(99.04%)
6 layers 61.39% 22.84% 46.54% 83.31% 61.77% 61.08% 68.32% 56.19% 33.2% 65.69% 56.04%(97.87%)
4 layers 63.28% 21.01% 43.47% 83.14% 60.82% 60.48% 67.97% 54.48% 33.8% 64.08% 55.25%(96.51%)
2 layers 54.54% 19.10% 41.90% 81.47% 57.22% 57.38% 65.84% 53.27% 33.7% 63.19% 52.76%(92.16%)
1 layer 53.16% 16.96% 33.29% 81.20% 51.89% 55.83% 64.50% 45.51% 30.1% 57.64% 49.01%(85.61%)
@@
40 layers 67.60% 25.19% 50.26% 83.61% 68.38% 63.29% 71.64% 60.21% 37.2% 68.70% 59.61%
10 layers 65.17% 23.56% 48.27% 83.57% 66.58% 62.01% 70.75% 59.13% 36.9% 67.39% 58.33%(97.85%)
9 layers 66.47% 23.65% 49.29% 83.74% 65.61% 62.31% 72.14% 59.71% 37.7% 67.29% 58.80%(98.64%)
Bunny-Phi3-mini-4B-V
32 layers 63.62% 29.19% 48.07% 83.69% 72.94% 62.35% 76.75% 60.64% 42.4% 72.09% 61.17%
8 layers 61.96% 27.21% 46.95% 83.11% 71.74% 61.38% 75.71% 59.69% 42.3% 71.53% 60.16%(98.35%)
Table 3: Performance of LVLMs with varying LLM backbones and parameter scales tuned with different numbers of layers. Values in parentheses denotes the percentage relative to the performance achieved by tuning all layers.
@@ 5.1 Generalizability Validation
To validate our findings of the visual region beyond Bunny-Llama-3-8B-V, we take LLaVA-1.5-7B, LLaVA-1.5-13B and Bunny-Phi3-mini-4B-V as additional testbeds to assess the generalizability across LVLMs with different LLM backbones and parameter scales. Following the setup in Sec. 4.2, we re-train these models with different number of layers that are sparsely and uniformly distributed within their respective backbones, including Vicuna-1.5-7B, Vicuna-1.5-13B and Phi-3-mini-4B Abdin et al. (2024). Results presented in Table 3 show that under our visual region positioning strategy, tuning approximately 25% of the layers consistently yield 98% of the full performance. This demonstrates that our approach generalizes effectively across varying LVLMs.
@@ 5.2 Computational Cost
To demonstrate the efficiency of visual region-based tuning, we report the computational costs associated with tuning different numbers of layers across various models using the LoRA strategy. For fair comparison across setups with different numbers of GPUs (specifically A800 GPUs in this analysis), we compute the product of the number of GPUs and running hours as a measure of computational cost. From Figure 3, Table 2 and Table 3, tuning a visual region comprising up to 25% of layers (8 layers for LLaVA-1.5-7B and Bunny-Llama3-8B-V, 10 layers for LLaVA-1.5-13B) can achieve 98% of full performance while achieving substantial reductions in computational overhead. Specifically, we reduce training time by 23% for LLaVA models and 13% for Bunny. These results highlight that the effectiveness of visual region-based tuning in training LVLMs efficiently with minimal performance trade-offs. Moreover, this relative reduction in computational cost would be more significant as dataset and model sizes scale.
@@ 5.3 Evaluation of Textual Tasks
As highlighted in Dai et al. (2024); Agrawal et al. (2024) and illustrated in Figure 1, multimodal training risks degradation of LLMs’ pre-trained linguistic knowledge and reasoning capabilities. To verify whether training our sparsely and uniformly distributed visual region affects the model linguistic capacity, we extend our evaluation to four text-only question answering datasets, MMLU Hendrycks et al. (2020), C-Eval Huang et al. (2023), CMMLU Li et al. (2023b), and BIG-bench-Hard Suzgun et al. (2022), covering diverse topics and fields. We use “Answer with the option’s letter from the given choices directly” as the prompts for the first three and “Please answer this question in a word or phrase” for BIG-bench-Hard, and allow models to provide explanations alongside its responses. We adopt a five-shot prompting strategy for MMLU, C-Eval and CMMLU, and a zero-shot strategy for BIG-bench-Hard.
Model Version MMLU BIG-Bench-H C-Eval CMMLU
@@
Fully-trained (32layers) 60.27% 30.93% 45.84% 45.68%
Partial-trained (8layers) 63.36% 31.50% 49.70% 48.39%
LLM-Backbone 66.01% 57.93% 50.52% 50.70%
@@
Fully-trained (32layers) 50.52% 26.85% 38.34% 37.27%
Partial-trained (8layers) 50.74% 31.64% 39.08% 37.71%
LLM-Backbone 49.78% 29.33% 38.78% 36.60%
Table 4: Performance on text-only tasks. The LLm backbones of Bunny-LLaMA3-8B-V and LLaVA-1.5-7B are respectively LLaMA3-8B and Vicuna-1.5-7B.
As shown in Table 4, fully-trained LVLMs generally exhibit decreased performance on text-only tasks compared to their LLM backbones, particularly with more powerful LLaMA3-8B and on the challenging BIG-bench-Hard dataset. In contrast, our selectively trained LVLMs minimally compromise models’ linguistic capacity, which consistently outperform fully-trained LVLMs, and sometimes even surpass their LLMs backbones. These results support our hypothesis that positioning the visual region strategically by tuning sparsely and uniformly distributed layers better preserves LLMs’ linguistic knowledge and reasoning capabilities, whereas full training may cause minor disruptions.
@@ 6 Visual Region-Based Layer Pruning
Building on our visual region targeted training, we propose a visual region-based pruning paradigm that selectively prunes less-important layers outside the visual region after training. Specifically, we follow the angular distance based layer importance metric and select 0$\sim$4 layers with the lowest angular distance outside the visual region. We do not evaluate pruning beyond this range as removing additional layers in LVLMs would lead to significant performance collapse. We evaluate this approach on LLaVA-1.5-7B across four datasets: OCRVQA, TextVQA, DocVQA and SciQA. As shown in Figure 4, our paradigm generally maintain higher performance, especially when pruning 3$\sim$4 layers, even though the visual region targeted trained model performs slightly worse than fully trained model without pruning. This result demonstrates that our paradigm effectively minimizes performance degradation compared to pruning in full-layer trained LVLMs, serving as an initial exploration into LVLM-specific pruning strategies.
@@ 7.1 Efficient Training and Inference
Recent research community has witnessed an emergent interest in LLMs Touvron et al. (2023); Chiang et al. (2023) and LVLMs Li et al. (2023c); Zhu et al. (2023); Bai et al. (2023); Liu et al. (2024) due to their remarkable ability to interpret and interact with the world via linguistic and visual channels. With the sustainably increased scale of LLMs and LVLMs, training or inference using all model parameters are cost for practical deployment. There are numerous techniques for efficient model training and inference. For instance, quantization reduce the memory footprint of models by decreasing the precision of model weights Dettmers et al. ; Dettmers and Zettlemoyer (2023); Xiao et al. (2023). Low rank adapters enable cost-effective fine-tuning by updating only a small subset of the adapter parameters Hu et al. (2021); Karimi Mahabadi et al. (2021).
@@ 8 Conclusion
Our work primarily focuses on LLaVA-1.5 family, Bunny-LLama3-8B-V and Bunny-Phi3-mini-4B-V to demonstrate the effectiveness and efficiency of our proposed training and inference paradigms for LVLMs. Future work will expand to a broader range of models to further validate the generalizability of our approach. Additionally, we will explore extensions to other modalities such as speech, and investigate the existence of other modality-specific regions to develop more versatile and scalable multimodal models.